# Quantum Multimodal Contrastive Learning Framework

### v3

@@ 2.7.4 QSTConv-GA
2.7.5 Quantum Encoding Layer
The quantum encoding layer is to integrate quantum computing capabilities into classical neural networks. By incorporating a quantum circuit within a classical deep learning model, the quantum encoding layer aims to leverage the unique properties of quantum mechanics, such as superposition and entanglement, to enhance the performance and capabilities of machine learning models. It leverages a parameterized quantum circuit defined by mapping classical EEG features as a set of qubits $\mathbf{x}\in\mathbb{R}^{n_{\text{qubits}}}$ and has trainable weights $\mathbf{w}\in\mathbb{R}^{n_{\text{layers}}\times n_{\text{qubits}}}$. The first step in the quantum encoding layer is encoding the input. Each qubit $q_{i}$ is rotated based on the input data $x_{i}$ using the rotation gate RY (rotation-Y gate):
$RY(x_{i})=\exp\left(-i\frac{x_{i}}{2}\sigma_{y}\right),$ (1)
where $\sigma_{y}$ is the Pauli-Y matrix:
$\sigma_{y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}.$ (2)
Second, there are some parameterized layers in the quantum encoding layer module, in this QEEGNet model, we use a ring pattern of CNOT (controlled-NOT) gates. The CNOT gate is a two-qubit gate where the state of one qubit (the control qubit) determines whether to flip the state of another qubit (the target qubit). The CNOT gate is a two-qubit quantum gate represented by the following 4x4 unitary matrix:
${CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$ (3)
A ring pattern involves applying CNOT gates such that each qubit is entangled with its neighbor, and the last qubit is entangled with the first qubit, forming a closed loop or "ring." This pattern ensures that entanglement is spread throughout the entire set of qubits. For each layer $l$ in $n_{\text{layers}}$, a ring pattern of CNOT gates entangles the qubits:
$CNOT(q_{i},q_{(i+1)\mod n_{\text{qubits}}})\quad\text{for}\quad i=0,1,\ldots,n_{\text{qubits}}-1.$ (4)
Each qubit $q_{i}$ undergoes an additional rotation based on the trainable weight $w_{l,i}$:
$RY(w_{l,i})=\exp\left(-i\frac{w_{l,i}}{2}\sigma_{y}\right).$ (5)
The final stage of the quantum encoding layer module is measurement the states of the qubits. The expectation value of the Pauli-Z operator $\sigma_{z}$ is measured for each qubit, yielding the output:
where $U$ is the unitary operation representing the quantum circuit. The Pauli-Z operator, also known as the Pauli-Z matrix, is represented by the following 2x2 matrix:
$\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix}.$ (7)
This operator is used to measure the z-component of the spin of a qubit and has eigenvalues of +1 and -1, corresponding to the basis states $|0\rangle$ and $|1\rangle$, respectively. The quantum encoding layer integrates quantum computing capabilities into classical neural networks. By embedding a quantum circuit within a classical deep learning model, the quantum encoding layer leverages the unique properties of quantum mechanics—such as superposition and entanglement—to enhance the performance and capabilities of machine learning models. This layer utilizes a parameterized quantum circuit, where classical EEG features are mapped to a set of qubits $\mathbf{x}\in\mathbb{R}^{n_{\text{qubits}}}$, with trainable weights $\mathbf{w}\in\mathbb{R}^{n_{\text{layers}}\times n_{\text{qubits}}}$.
@@ 2.7.5 Quantum Encoding Layer
where $\sigma_{y}$ is the Pauli-Y matrix:
$\sigma_{y}=\begin{pmatrix}0&-i\\ i&0\end{pmatrix}.$ (9)
2. Parameterized Layers: The quantum encoding layer uses a ring pattern of controlled-NOT (CNOT) gates. A CNOT gate, represented by the following $4\times 4$ unitary matrix, flips the state of a target qubit based on the state of a control qubit:
$\text{CNOT}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&0&1\\ 0&0&1&0\end{pmatrix}.$ (10)
The ring pattern involves applying CNOT gates to entangle each qubit with its neighbor, forming a closed loop. For each layer $l$ in $n_{\text{layers}}$, the CNOT gates are applied as follows:
@@
where $U$ represents the unitary operation of the quantum circuit. The Pauli-Z operator, represented by the matrix:
$\sigma_{z}=\begin{pmatrix}1&0\\ 0&-1\end{pmatrix},$ (14)
is used to measure the z-component of the qubit’s spin, with eigenvalues $+1$ and $-1$ corresponding to the basis states $|0\rangle$ and $|1\rangle$, respectively.
2.7.6 Quantum Multimodal Contrastive Learning
@@
3 Result and Conclusion
3.1 Datasets, Preprocessing and Experiment Details
@@
3.2 Result and Conclusion
The experiment results are shown in TABLE 1 and TABLE 2. The experimental results demonstrate that the QNervFormer-GA model outperforms the other models in both top-1 and top-5 validation accuracy across the 10 subjects in the ThingsEEG dataset. Specifically, QNervFormer-GA achieved the highest average top-1 accuracy at 5.2%, closely followed by QNervFormer at 5.1%. QSTConv-GA and QSTConv lagged behind with average top-1 accuracies of 4.0% and 3.5%, respectively. In terms of top-5 accuracy, QNervFormer-GA again led with an average of 20.3%, slightly ahead of QNervFormer at 20.2%. QSTConv-GA recorded an average of 16.8%, while QSTConv was the lowest at 13.7%. These results highlight the superior generalization ability of QNervFormer-GA, making it the most effective model among those tested for EEG-based image recognition on the ThingsEEG dataset.

### v4

@@ III-B4 QSTConv-GA
III-C Quantum Encoding Layer
The quantum encoding layer integrates quantum computing capabilities into classical neural networks. By embedding a quantum circuit within a classical deep learning model, the quantum encoding layer leverages the unique properties of quantum mechanics—such as superposition and entanglement—to enhance the performance and capabilities of machine learning models. This layer utilizes a parameterized quantum circuit, where classical EEG features are mapped to a set of qubits $\mathbf{x}\in\mathbb{R}^{n_{\text{qubits}}}$, with trainable weights $\mathbf{w}\in\mathbb{R}^{n_{\text{layers}}\times n_{\text{qubits}}}$.
@@ III-C Quantum Encoding Layer
where $\sigma_{y}$ is the Pauli-Y matrix.
2. Parameterized Layers: The quantum encoding layer uses a ring pattern of controlled-NOT (CNOT) gates. The ring pattern involves applying CNOT gates to entangle each qubit with its neighbor, forming a closed loop. For each layer $l$ in $n_{\text{layers}}$, the CNOT gates are applied as follows:
@@
where $U$ represents the unitary operation of the quantum circuit.
III-D Quantum Multimodal Contrastive Learning
@@
IV Result and Conclusion
IV-A Datasets, Preprocessing and Experiment Details
@@
IV-B Result and Conclusion
The experiment results are shown in TABLE I and TABLE II. The experimental results demonstrate that the QNervFormer-GA model outperforms the other models in both top-1 and top-5 validation accuracy across the 10 subjects in the ThingsEEG dataset. Specifically, QNervFormer-GA achieved the highest average top-1 accuracy at 5.2%, closely followed by QNervFormer at 5.1%. QSTConv-GA and QSTConv lagged behind with average top-1 accuracies of 4.0% and 3.5%, respectively. In terms of top-5 accuracy, QNervFormer-GA again led with an average of 20.3%, slightly ahead of QNervFormer at 20.2%. QSTConv-GA recorded an average of 16.8%, while QSTConv was the lowest at 13.7%. These results highlight the superior generalization ability of QNervFormer-GA, making it the most effective model among those tested for EEG-based image recognition on the ThingsEEG dataset.